OpenStudy (anonymous):

an airplane maintains a speed of 567 km/h relative to the air it is flying through as it makes a tripto a city 780 km away to the north a) what time interval is rquired for the trip if the plane flies through a headwind blowing at 33.0 km/h toawrd the south? b) what time interval is required if theres a tailwind with same speed? c) what time interval required if there is a crosswind blowing at 33 km/h to the east relative to the ground?

OpenStudy (dumbcow):

a) it is flying into the wind, so speed is 567 -33 = 534 d = r*t t = d/r = 780/534 = 1.46 hr b) now flying with the wind so add 33. 567 + 33 = 600 t = 780/600 = 1.3 hr c) see drawing x = sqrt(567^2 -33^2) = 566.04 t = 780/566.04 = 1.378 hr